3. A marble of mass 10 g falling through a tank of liquid experiences a resistive force R. proportional to its velocity. Specifically, R =-bv, where b = 3.5 × 10-2 kg s-1. The marble is dropped from rest and begins to accelerate downward (a) What is the marbles terminal velocity in this liquid? (b) How mich time elapses before the marble has reached 95% of its terminal velocity? (c) How far does it travel during this time?
Please make work legible and clear I want to understand how to do it. Please also only use Newtonian mechanics
Answer #1

Applying Newton's second law

F_{N}=m\cdot a

but the net force acting on the marble is

F_{N}=W-b\cdot v=m\cdot g-b\cdot v

where:
R=-b\cdot v  is the resistive force

W=m\cdot g  is the weight of the marble

combining equations

m\cdot a=m\cdot g-b\cdot v

but remember that a=\frac{dv}{dt}

replacing in equation

m\cdot \frac{dv}{dt}=m\cdot g-b\cdot v  dividing by m \frac{dv}{dt}= g-\frac{b}{m}\cdot v

part A.

speed is terminal when the acceleration is zero, ie.a= \frac{dv}{dt}=0

Expression of terminal velocity 0= g-\frac{b}{m}\cdot v_{T}\Rightarrow v_{T}=\frac{g\cdot m}{b}

where:

the mass is m=10g\cdot \frac{1kg}{10^{3}g }=10\cdot 10^{-3}kg

the constant b=3.5\cdot 10^{-2}kg\cdot s^{-1}

acceleration of gravity g=9.81\frac{m}{s^{2}}

evaluate numerically

v_{T}=\frac{g\cdot m}{b}=\frac{9.81\frac{m}{s^{2}}\cdot 10\cdot 10^{-3}kg}{ 3.5\cdot 10^{-2}kg\cdot s^{-1}}=2.80\frac{m}{s}

The terminal velocity is

v_{T}=2.80\frac{m}{s}

Part B

The expression that satisfies the equation \frac{dv}{dt}= g-\frac{b}{m}\cdot v   

considering
initial velocity v_{i}=0\frac{m}{s}
initial time t_{i}=0s

V=V_{T}\left ( 1-e^{-\frac{t }{\tau }} \right )

where:

terminal velocity

v_{T}=2.80\frac{m}{s}

time constant

\tau =\frac{ m}{b }=\frac{ 10\cdot 10^{-3}kg}{ 3.5\cdot 10^{-2}kg\cdot s^{-1}}=0.29s

if you want to know when

V=0.95V_{T}

substituting in the equation

0.95\cdot V_{T}=V_{T}\left ( 1-e^{-\frac{t }{\tau }} \right )

simplifying

0.95= 1-e^{-\frac{t }{\tau }}

expression for t

0.95= 1-e^{-\frac{t }{\tau }} \Rightarrow e^{-\frac{t }{\tau }}=1-0.95\Rightarrow e^{-\frac{t }{\tau }}=0.05

applying logarithmic properties

ln\left (e^{-\frac{t }{\tau }} \right )=ln\left (0.05 \right )\Rightarrow -\frac{t }{\tau }=ln\left (0.05 \right )\Rightarrow t=-\tau \cdot ln\left ( 0.05 \right )

evaluate numerically t=-\tau \cdot ln\left ( 0.05 \right )=-0.29s\cdot ln\left ( 0.05 \right )=0.87s

finally t=0.87s

Part C

the position of the object is equal to v=\frac{dx}{dt}

you can write the equation V=V_{T}\left ( 1-e^{-\frac{t }{\tau }} \right ) how \frac{dx}{dt}=V_{T}\left ( 1-e^{-\frac{t }{\tau }} \right )

organizing

dx=V_{T}\left ( 1-e^{-\frac{t }{\tau }} \right )dt

The position is equal to

\int_{x_{0}}^{x}dx=V_{T}\int_{t_{0}}^{t}\left ( 1-e^{-\frac{t }{\tau }} \right )dt

organizing integrals

\int_{x_{0}}^{x}dx=V_{T}\int_{t_{0}}^{t} 1dt-V_{T}\int_{t_{0}}^{t}e^{-\frac{t }{\tau }}dt

solving each integral

\left [x \right ]_{x_{0}}^{x}=\left [V_{T}\cdot t+V_{T}\cdot \tau \cdot e^{-\frac{t }{\tau }} \right ]_{t_{0}}^{t}  considering x_{0}=0  and t_{0}=0

position equation

x=V_{T}\cdot t+V_{T}\cdot \tau \cdot e^{-\frac{t }{\tau }}

evaluating in t=0.87s

x=V_{T}\cdot t+V_{T}\cdot \tau \cdot e^{-\frac{t }{\tau }}=2.80\frac{m}{s}\cdot 0.87s+2.80\frac{m}{s}\cdot 0.29s\cdot e^{-\frac{0.87 }{ 0.29s}}=2.44m\dotplus 16.31m=18.75m

x=18.75m